A farm owner is considering replacing his obsolete tractor with one of two new state-of-the- tractors. This new machine would cost $125,000 and would have a ten-year useful life Unfortunately, the new machine would have no salvage value but would result in annual cost savings of $23,000 per year. The current old tractor can be sold now for S10,000. The farm owner's Cost of Capital is 10%. The farm owner uses the straight line method of depreciation (this depreciation information is needed only for calculating the "Simple Rate of Return" in Question #3). 


 a.) Calculate the Net Present Value of replacing the raster. 

b.) Based on this method of comparison, would you recommend replacing the tractor? Why?
Answer #1

a. Net present value

1. Net Investment

Cost of new tractor. = 125,000$

Sale proceeds of old tractor = (10,000$)

Net investment. = 115,000$

2. Present value of cash inflows

Annual savings = 23,000$

Present value of savings = 23,000 PVAF10%10years = 23,000 * 6.1446 = 141,326

3. Net present value = Cash inflows - net investment = 141,326 - 115,000 = 26,326$

b. Replacement Decision

It is advised to replace the old tractor with new one since the net present value is positive.

​​